# Invariant measures for some dissipative systems from the Jacobi last multiplier

### v2

@@ 5 Discussion
In this paper, we have presented invariant phase-space measures for certain mechanical systems using the framework of Jacobi last multipliers. In particular, we analyzed conformal Hamiltonian dynamics as well as contact Hamiltonian dynamics. Another class of systems that we analyzed were systems of the Liénard class which may possess position-dependent damping. For such systems, we discussed a geometric description of the dynamics on cotangent bundles by presenting the notion of a generalized conformal vector field. This also allowed us to obtain analytical expressions for invariant phase-space measures associated with such systems when the Cheillini integrability condition is satisfied. A closely-related analysis may be carried out for a class of Levinson-Smith equations discussed recently in [23].
@@
where we have taken the modulus sign in writing $|M|$ as is necessary for interpreting it as a phase-space density. The above-mentioned equation (5.1) may be straightforwardly interpreted as a generalized Liouville equation (see also, [27]) where $|M|$ can be viewed as the phase-space density. For conservative mechanical systems described by Hamiltonian dynamics on symplectic manifolds (where $X=X_{H}$; as defined in (2.6)), from £$\pounds_{X_{H}}\omega^{n}=0$, one has ${\rm div}X_{H}=0$ which gives $X_{H}(\ln|M|)=0=X_{H}(|M|)$. This is the Liouville equation that one encounters in classical mechanics and statistical mechanics, wherein one identifies the phase-space density $\rho\equiv|M|$. Explicitly, one has
@@
which is foundational to the formulation of statistical mechanics. In the present study, we have considered systems that cannot be described from the point of view of Hamiltonian dynamics on symplectic manifolds and have presented explicit results for invariant phase-space measures $\{\Omega_{\rm Invariant}=M\Omega\}$ starting with volume-forms $\{\Omega\}$ that are not conserved under the dynamics, i.e., £$\pounds_{X}\Omega\neq 0$. Different cases for $X$ as discussed in this work are conformal ($X^{\gamma}_{H}$), generalized conformal ($X^{K}_{H}$), and contact ($X_{h}$) vector fields. Thus, the present exposition can be understood as a preliminary step towards a formulation of statistical mechanics for dissipative systems (see for example, [11, 13]).

### v3

@@ 5 Discussion
In this paper, we have presented invariant phase-space measures for certain mechanical systems using the framework of Jacobi last multipliers. In particular, we analyzed conformal Hamiltonian dynamics as well as contact Hamiltonian dynamics. Another class of systems that we analyzed were systems of the Liénard class which may possess position-dependent damping. For such systems, we discussed a geometric description of the dynamics on cotangent bundles by presenting the notion of a generalized conformal vector field. This also allowed us to obtain analytical expressions for invariant phase-space measures associated with such systems when the Chiellini integrability condition is satisfied. A closely-related analysis may be carried out for a class of Levinson-Smith equations discussed recently in [23].
@@
where we have taken the modulus sign in writing $|M|$ as is necessary for interpreting it as a phase-space density. The above-mentioned equation (5.1) may be straightforwardly interpreted as a generalized Liouville equation (see also, [27]) where $|M|$ can be viewed as the phase-space density. For conservative mechanical systems described by Hamiltonian dynamics on symplectic manifolds (where $X=X_{H}$; as defined in (2.6)), from $\pounds_{X_{H}}\omega^{n}=0$, one has ${\rm div}X_{H}=0$ which gives $X_{H}(\ln|M|)=0=X_{H}(|M|)$. This is the Liouville equation that one encounters in classical mechanics and statistical mechanics, wherein one identifies the phase-space density $\rho=|M|$. Explicitly, one has
@@
which is foundational to the formulation of statistical mechanics. In the present study, we have considered systems that cannot be described from the point of view of Hamiltonian dynamics on symplectic manifolds and have presented explicit results for invariant phase-space measures $\{\Omega_{\rm Invariant}=M\Omega\}$ starting with volume-forms $\{\Omega\}$ that are not conserved under the dynamics, i.e., $\pounds_{X}\Omega\neq 0$. Different cases for $X$ as discussed in this work are conformal ($X^{\gamma}_{H}$), generalized conformal ($X^{K}_{H}$), and contact ($X_{h}$) vector fields. Thus, the present exposition can be understood as a preliminary step towards a formulation of statistical mechanics for dissipative systems (see for example, [11, 13]).